OpenStudy (anonymous):

Which equation is an identity? A) 11-(2v+3)=-2v-8 B) 5w+8-w=6w-2(w-4) C) 7m-2=8m+4-m D) 8y+9=8y-3 I think the answer is D, but correct me if I am wrong.

OpenStudy (mertsj):

B

OpenStudy (mertsj):

An identity means they are the same thing for all values of the variable.

OpenStudy (haseeb96):

Answer is B because 5w+8-w=6w-2(w-4) 4w+8=6w-2w+8 4w+8=4w+8 both sides has 4w+8 so we will cancel it and the answer is 1 in general we say it identity